Deduce the structure of organic compound based on data given below :
Mol. Formula : C9H10O2
M/S : 150, 108, 91, 77, 65, 51 and 43.
1
Expert's answer
2019-11-11T07:47:32-0500

C6H3(CH3)(OH)(CH2-CH=O) - 150.

-CH2-CH=O - 43.